OpenStudy (bbb911):

This is Geometry: Find the lateral area of a regular octagonal pyramid if the slant height is 12 meters and the length of each side of its base is 6 meters.

OpenStudy (amistre64):

lateral area is side area

OpenStudy (anonymous):

288

OpenStudy (amistre64):

slant height doesnt seem to be an agreed upon term; but I assume it means face height

OpenStudy (bbb911):

thanks him :D

OpenStudy (amistre64):

8 times the area of a face (12*6/2)

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

4*12*6 ------ = 4*6*6 = 4*36 = 120+24 = 144 right? 2

OpenStudy (bbb911):

no its 288

OpenStudy (amistre64):

yeah; I see my error; I mentally halved the 8 and forgot to account for it lol

OpenStudy (bbb911):

hehe its ok

OpenStudy (amistre64):

\[\frac{8*12*36}{2}=288\]

OpenStudy (amistre64):

.... did it again lol

OpenStudy (anonymous):

growing old amistre

OpenStudy (amistre64):

growing? lol... Ive been old...

OpenStudy (anonymous):

the sachin tendulkar of openstudy